how much pure gold should be melted with 15 grams of 14-karat gold to produce 18-karat gold?
Answer by ankor@dixie-net.com(22740)

how much pure gold should be melted with 15 grams of 14-karat gold to produce 18-karat gold?
:
Pure gold is 24 karat
:
Let x = amt of pure gold
:
A karat equation:
14(15) + 24(x) = 18(x + 15)
210 + 24x = 18x + 270
24x - 18x = 270 - 210
6x = 60
x = 10 grams of pure gold required
:
:
Check solution of x = 10:
14(15) + 24(10) = 18(25)